OpenStudy (anonymous):

show steps factor (m^5+5m^4-2m^3-10m^2+m+5)

OpenStudy (anonymous):

m^4(m+5)-2m^2(m+5)+(m+5) =(m+5)(m^4-2m^2+1) =(m+5)(m^2-1)^2 \[=(m+5)(m+1)^2(m-1)^2\]